# PixelCAM: Pixel Class Activation Mapping for Histology Image Classification and ROI Localization

### v1

@@ 4 Results and Discussion
4.1 Experimental Methodology
(a) Datasets. Our experiments are performed on two standard public datasets for WSOL in histology images. In particular, we use GlaS dataset [24] for colon cancer, and the patch version of CAMELYON16 [10, 22] for breast cancer. In both cases, we follow the same WSOL experimental protocol used in [3, 12, 22].
GlaS.
The GlaS dataset is used for the diagnosis of colon cancer. The dataset consists of 165 images from 16 Hematoxylin and Eosin (H&E) and includes labels at both pixel-level and image-level (benign or malign). The dataset consists of 67 images for training, 18 for validation, and 80 for testing. We use the same protocol as in [22, 12, 3]. Similarly, we use 3 samples per class with full supervision in validation set for model selection for localization (B-LOC).
CAMELYON16.
A patch-based benchmark [22] is extracted from the CAMELYON16 dataset that contains 399 Whole slide images categorized into two classes (normal and metastasic) for the detection of breast cancer metastases in H&E-stained tissue sections of sentinel lymph nodes. Patch extraction of size ${512\times 512}$ follows a protocol established by [22, 12, 3] to obtain patches with annotations at the image and pixel level. The dataset contains a total of 48870 images, including 24348 for training, 8850 for validation, and 15664 for testing. From the validation dataset, 6 examples per class are randomly selected to be fully supervised to perform model selection for localization (B-LOC) similarly to [22, 12, 3].
(b) Implementation details. To assess the impact of our method, we first train a baseline WSOL method alone, and then contrast its performance when combined with our method. The pseudo-labels are extracted from the baseline model. Different recent WSOL baseline methods are used from single- and tow-step families, including DeepMIL [13], GradCAM++ [6], LayerCAM [14], SAT [27], and NEGEV [3]. For CNN-based models, we used ResNet-50 as the backbone, and for SAT, the transformer-based model DeiT-Tiny is used. In terms of performance measures, standard WSOL metrics are used, including image classification accuracy, and pixel-localization accuracy PxAP [7, 5, 3, 12]. In addition, pixel-wise true/false positive/negative rates are used as well. For the localization task, we compare all WSOL methods to the full-supervised case using the U-Net model.
For pretraining a WSOL baseline method on the data, we use the same setup as in [12]. The first part of the model training is defined using the backbone trained on ImageNet [16]. The training is done using SGD with a batch size of 32 [12]. For the GlaS dataset, training is performed over 1000 epochs, and 20 epochs for CAMELYON16. A weight decay of $10^{-4}$ is also used. During training, images are resized to ${256\times 256}$, then randomly cropped to ${224\times 224}$. A hyperparameter search was conducted for the learning rate parameters among the values {0.0001, 0.001, 0.01}, and its decaying factor among {0.1, 0.4, 0.9} following [12]. In the second phase of training PixelCAM, we use the CAMs generated by the previous method and continue the training using the same setup as defined previously. For the OOD scenario, the source model is trained using the same setup on a source dataset and evaluated on a target dataset.
@@ 4.1 Experimental Methodology
DeepMIL [13] $\dagger$ icml 79.9 100.0 71.3 85.0
DeepMIL w/ NEGEV [3] $\star$ midl 85.9 100.0 75.6 89.9
@@
GradCAM++ [6] $\dagger$ wacv 76.8 100.0 49.1 63.4
GradCAM++ w/ NEGEV [3] $\star$ midl 77.1 100.0 68.6 89.4
@@
LayerCAM [14] $\dagger$ ieee tip 75.1 100.0 33.2 84.8
LayerCAM w/ NEGEV[3] $\star$ midl 73.8 100.0 66.8 89.1
@@ 4.2 WSOL Results
We first evaluate PixelCAM using the standard WSOL protocol over GlaS and CAMELYON16 dataset [22]. Results are reported in Tab. 1. Combining our method with a WSOL baseline often provides an improvement in localization accuracy. Over GlaS dataset, our method improves the PxAP performance by (+5.6%,+9.8%,+8.5%) compared to the WSOL baseline methods (DeepMIL, GradCAM++, LayerCAM) respectively, while on CAMELYON16 dataset, our method gains +4.4%,+15.0%,+33.0%, respectively. Fig. 2 shows visual prediction examples. Additionally, our method improves classification performance as well. This indicates that well-separated classes at pixel-feature level in our model help improve localization but also facilitate global image classification. We further provide measures of how well are separated FG and BG pixel-features in Fig 3 on the GlaS test set. These results show that PixelCAM histogram of class separability index is shifted to the right compared to the baseline. This indicates a better class separability.
@@
Figure 3: Histogram of foreground/background separability index for WSOL baseline methods alone vs when combined with PixelCAM on GlaS test set. The higher the value, the better the separability is.
4.3 Class Separability Index
@@
In this section, we present the class separability index between foreground (FG) and background (BG) classes at pixel-feature level over test set of both datasets GlaS and CAMELYON16. To measure how well FG/BG classes are separated in the feature space of pixels, we resort to the class separability index [8]. The class separability, $J$, is based on the Within-class scatter matrix ($S_{W}$) and the Between-class scatter matrix ($S_{B}$) of pixel-features, defined as follows,
$\mathbf{S}_{W}=\sum_{i=1}^{c}\left[\sum_{j=1}^{n_{i}}(\mathbf{x}_{i,j}-\mathbf{m}_{i})(\mathbf{x}_{i,j}-\mathbf{m}_{i})^{\top}\right]\,,$ (4)
$\mathbf{S}_{B}=\sum_{i=1}^{c}n_{i}(\mathbf{m}_{i}-\mathbf{m})(\mathbf{m}_{i}-\mathbf{m})^{\top}\,,$ (5)
@@
where $c$ is the number of class (in our case ${c=2}$ for foreground and background). Note that this index $J$ is computed over a single image using the pixel-labels. Let $n_{i}(i=0,..,c)$ the number of pixels in the $i$-th class. The feature vector $\mathbf{x}_{i,j}\in\mathbb{R}^{d}$ denotes the $j$-th pixel of the $i$-th class. It represents the vector ${\bm{F}_{p}}$ for the pixel $p$ in the main paper. The mean vector of the $i$-th class is given by $\mathbf{m_{i}}$, while $\mathbf{x}$ represents the mean vector computed over all feature vectors. The class separability denoted $\mathbf{J}$ is defined as follows [8],
$\mathbf{J}=\frac{\mathbf{S}_{B}}{\mathbf{S}_{W}}\,.$ (6)
@@
In Table 2 and 3, we provide the average class separability over normal and cancer classes and the overview over the entire dataset. As we can observe, PixelCAM improves the class separability between FG and BG of WSOL baseline methods on both dataset except for GradCAM++ on CAMELYON16. This separability is explained by the high number of images with a low separability as illustrated in Fig 4. The improvement of PixelCAM in the OOD scenario compared to the WSOL baseline methods such as DeepMIL and LayerCAM on GlaS and CAMELYON16 can be attributed to better separability as shown in Figs 5 and 6. We provide examples of features separability at the pixel level in section E.
@@
4.4 Domain Shift Results
CAMELYON16 $\rightarrow$ GlaS GlaS $\rightarrow$ CAMELYON16
WSOL models Venue PxAP $\uparrow$ CL $\uparrow$ PxAP $\uparrow$ CL $\uparrow$
DeepMIL [13] icml 64.5 81.2 29.0 55.2
DeepMIL w/ PixelCAM - 69.1 83.8 30.2 52.5
GradCAM++ [6] wacv 52.9 53.7 39.1 52.4
GradCAM++ w/ PixelCAM - 56.2 71.2 36.9 63.3
LayerCAM [14] ieee tip 59.5 77.5 30.4 51.4
LayerCAM w/ PixelCAM - 67.2 78.8 34.8 55.8
SAT [27] iccv 50.7 67.5 24.3 50.2
SAT w/ PixelCAM - 55.1 78.8 22.5 50.4
Table 4: OOD results: Localization (PxAP) and classification (CL) accuracy on target test dataset: GlaS and CAMELYON16.
Additional experiments were conducted to assess the ability of PixelCAM to train robustness models on OOD data. To this end, consider the domain shift between GlaS (colon cancer) and CAMELYON16 (breast cancer)This simulates a domain adaptation problem in the source-only case, where a model is pre-trained on source domain data, and then evaluated on target data. Results are reported in Tab. 4. As shown in [12], WSOL methods typically decline when facing OOD in histology data. Overall, our PixelCAM method yields better performance over both tasks and across all baselines. This is mainly due to better separated pixel-features learned by our model which make them more robust to OOD data. This is demonstrated in 4.3 by inspecting the class separability of pixel-features of the target data. As presented in Tab. 4, using PixelCAM outperforms in general WSOL baselines alone on the case CAMELYON16 $\rightarrow$ GlaS. This improves PxAP scores by 4.6%, 3.3%, 7.7%, and 4.4% compared to DeepMIL, GradCAM++, LayerCAM, and SAT, respectively. Moreover, PixelCAM also improves image classification performance. The other case, GlaS $\rightarrow$ CAMELYON16, is much more challenging since the train source dataset is very small, compared to the very large and difficult target set. While the overall performance is low compared to when GlaS is the target, our method still yields better localization performance in general. Moreover, classification performance is improved overall across both scenarios. More results and interpretations are included in Appendices  D and  E.
5 Conclusion
In this paper, we propose PixelCAM, a novel multi-task single-step WSOL method. It is based on a cost-effective FG/BG pixel-wise classifier in the pixel-feature space, allowing for spatial object localization. It leverages pixel-pseudo labels cues for localization. It aims at explicitly learning discriminant pixel-features and accurately separate FG/BG regions, promoting better ROI localization and image classification. Both tasks are optimized in parallel without incurring notable computational cost nor additional parameters. Our pixel classifier is versatile. It can easily be integrated into any CNN- or transform-based classifier architecture without modification. Results on two histology datasets showed that PixelCAM can improve WSOL baseline methods over both tasks, but also make them robust to OOD data. However, there is still room to improve our method by better optimizing both tasks while reducing their mutual negative impact. In addition, dealing with OOD in WSOL scenario is still an ongoing issue, and our method still requires more improvements despite its progress.
Acknowledgement
This appendix contains the following points:
  1. 1.
    More localization performance analysis
  2. 2.
    Computational complexity
  3. 3.
    Limitations and future work
  4. 4.
    Ablations:
    1. (a)
      Impact of PixelCAM depth
    2. (b)
      Impact of model selection of CAM pre-trained model
    3. (c)
      Impact of pixel sampling technique
    4. (d)
      Impact of the number of sampled pixels $n$
    5. (e)
      Impact of $\lambda$
    6. (f)
      Impact of using a pretrained WSOL model on same and external dataset for pseudo-labeling
  5. 5.
    More visualizations results:
    1. (a)
      Standard WSOL setup
    2. (b)
      Over Target set with OOD setup
Appendix A More Localization Performance Analysis
PxAP is the primary metric used to measure localization performance. However, following [22], we include other pixel-wise performance measures that is true/false positives/negative rates. As shown previously, our method PixelCAM achieves better results in terms of PxAP performance. This improvement is observed in Tab. 5 with a higher number of true positive/negative rates compared to GradCAM++, LayerCAM, SAT, and NEGEV on the GlaS dataset. For the CAMELYON16 dataset, we observe that PixelCAM increases the true positive rate compared to standard methods while being competitive to NEGEV. However, we note a significant improvement when applying PixelCAM to the SAT method, with a considerable increase in true positives, thereby reducing the over-activation issue.

### v2

@@ 3 Results and Discussion
3.1 Experimental Methodology [^3] A more detailed description of datasets and implementation are in Appendices  B and C , respectively.
(a) Datasets. Our experiments are performed on two standard public datasets for WSOL in histology images. In particular, we use GlaS dataset Sirinukunwattana et al. (2015) for colon cancer, and the patch version of CAMELYON16 Ehteshami Bejnordi et al. (2017); Rony et al. (2023) for breast cancer. In both cases, we follow the same WSOL experimental protocol used in Belharbi et al. (2022a); Guichemerre et al. (2024); Rony et al. (2023).
(b) Implementation details. To assess the impact of our method, we first train a baseline WSOL method alone, and then contrast its performance when combined with our method. The pseudo-labels are extracted from the baseline model. Different recent WSOL baseline methods are used from single- and tow-step families, including DeepMIL Ilse et al. (2018), GradCAM++ Chattopadhyay et al. (2018), LayerCAM Jiang et al. (2021), SAT Wu et al. (2023), and NEGEV Belharbi et al. (2022a). For CNN-based models, we used ResNet-50 as the backbone, and for SAT, the transformer-based model DeiT-Tiny is used. In terms of performance measures, standard WSOL metrics are used, including image classification accuracy, and pixel-localization accuracy PxAP Choe et al. (2020); Belharbi et al. (2022c, a); Guichemerre et al. (2024). In addition, pixel-wise true/false positive/negative rates are used as well. For the localization task, we compare all WSOL methods to the full-supervised case using the U-Net model.
@@
DeepMIL Ilse et al. (2018) $\dagger$ icml 79.9 100.0 71.3 85.0
DeepMIL w/ NEGEV Belharbi et al. (2022a) $\star$ midl 85.9 100.0 75.6 89.9
@@
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 76.8 100.0 49.1 63.4
GradCAM++ w/ NEGEV Belharbi et al. (2022a) $\star$ midl 77.1 100.0 68.6 89.4
@@
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 75.1 100.0 33.2 84.8
LayerCAM w/ NEGEVBelharbi et al. (2022a) $\star$ midl 73.8 100.0 66.8 89.1
@@
GlaS CAMELYON16
WSOL models Venue T-stat $\uparrow$ p-value $\downarrow$ T-stat $\uparrow$ p-value $\downarrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 10.8 $1.5\times 10^{-11}$ 4.8 $4.1\times 10^{-5}$
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 29.1 $1.8\times 10^{-22}$ 17.9 $7.4\times 10^{-17}$
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 12.7 $4.0\times 10^{-13}$ 26.5 $2.2\times 10^{-21}$
SAT Wu et al. (2023) $\dagger$ iccv 58.3 $8.9\times 10^{-31}$ 99.7 $2.8\times 10^{-37}$
Table 2: T-test statistics between baselines (DeepMIL,GradCAM++,LayerCAM,SAT) and our method (PixelCAM) for localization performance.
@@ 3.2 WSOL Results
We first evaluate PixelCAM using the standard WSOL protocol over GlaS and CAMELYON16 dataset Rony et al. (2023). Results are reported in Tab. 1. Combining our method with a WSOL baseline often provides an improvement in localization accuracy. Over GlaS dataset, our method improves the PxAP performance by (+5.6%,+9.8%,+8.5%) compared to the WSOL baseline methods (DeepMIL, GradCAM++, LayerCAM) respectively, while on CAMELYON16 dataset, our method gains +4.4%,+15.0%,+33.0%, respectively. Fig. 2 shows visual prediction examples. The localization performance gains are supported by statistical significance, as confirmed by t-tests, in Tab. 2, yielding very low p-values (below 0.05). Additionally, our method improves classification performance as well. This indicates that well-separated classes at pixel-feature level in our model help improve localization but also facilitate global image classification. We further provide measures of how well are separated FG and BG pixel-features in Fig 3 on the GlaS test set. These results show that PixelCAM histogram of class separability index is shifted to the right compared to the baseline. This indicates a better class separability.
@@
Figure 3: Histogram of foreground/background separability index for WSOL baseline methods alone vs when combined with PixelCAM on GlaS test set. The higher the value, the better the separability is. The definition of class separability is in Appendix D.
3.3 Out-of-Distribution Results
CAMELYON16 $\rightarrow$ GlaS GlaS $\rightarrow$ CAMELYON16
WSOL models Venue PxAP $\uparrow$ CL $\uparrow$ PxAP $\uparrow$ CL $\uparrow$
DeepMIL Ilse et al. (2018) $\dagger$ icml 64.5 81.2 29.0 55.2
DeepMIL w/ NEGEVBelharbi et al. (2022a) $\star$ midl 62.7 81.2 27.8 52.9
DeepMIL w/ PixelCAM $\dagger$ - 69.1 83.8 30.2 52.5
GradCAM++ Chattopadhyay et al. (2018) $\dagger$ wacv 52.9 53.7 39.1 52.4
GradCAM++ w/ NEGEVBelharbi et al. (2022a) $\star$ midl 64.5 75.0 24.2 56.7
GradCAM++ w/ PixelCAM $\dagger$ - 56.2 71.2 36.9 63.3
LayerCAM Jiang et al. (2021) $\dagger$ ieee tip 59.5 77.5 30.4 51.4
LayerCAM w/ NEGEVBelharbi et al. (2022a) $\star$ midl 63.9 83.7 27.9 52.3
LayerCAM w/ PixelCAM $\dagger$ - 67.2 78.8 34.8 55.8
SAT Wu et al. (2023) $\dagger$ iccv 50.7 67.5 24.3 50.2
SAT w/ PixelCAM $\dagger$ - 55.1 78.8 22.5 50.4
Table 3: OOD results: Localization (PxAP) and classification (CL) accuracy on target test dataset: GlaS and CAMELYON16. The symnbol $\dagger$” refers to model with a one-step approach, while $\star$” refers to model from two-step family.
Additional experiments were conducted to assess the ability of PixelCAM to train robustness models on OOD data. To this end, consider the domain shift between GlaS (colon cancer) and CAMELYON16 (breast cancer). This simulates a domain adaptation problem in the source-only case, where a model is pre-trained on source domain data, and then evaluated on target data. Results are reported in Tab. 3. As shown in Guichemerre et al. (2024), WSOL methods typically decline when facing OOD in histology data. Overall, our PixelCAM method yields better performance over both tasks and across all baselines. This is mainly due to better separated pixel-features learned by our model which make them more robust to OOD data. This is demonstrated in the appendix D by inspecting the class separability of pixel-features of the target data. As presented in Tab. 3, using PixelCAM outperforms in general WSOL baselines alone on the case CAMELYON16 $\rightarrow$ GlaS. This improves PxAP scores by 4.6%, 3.3%, 7.7%, and 4.4% compared to DeepMIL, GradCAM++, LayerCAM, and SAT, respectively. Moreover, PixelCAM also improves image classification performance. The other case, GlaS $\rightarrow$ CAMELYON16, is much more challenging since the train source dataset is very small, compared to the very large and difficult target set. While the overall performance is low compared to when GlaS is the target, our method still yields better localization performance in general. Moreover, classification performance is improved overall across both scenarios. More results and interpretations are included in Appendices  J and  K.
We further evaluate the robustness of PixelCAM by altering the stainings in the test sets of GlaS using stain styles from the other dataset. The stainings were progressively modified by selecting stain variations ordered by their distance to the original stain. Stain 1 introduces a minor alteration, while stain 10 represents a substantial shift. This setup simulates an increasing staining shift while keeping the organ type unchanged. Additional results for CAMELYON16 are presented in Appendices H.
Figure 4: Classification (CL) accuracy on GlaS test sets with LayerCAM and PixelCAM with different stainings.
Figure 5: Localization (PxAP) accuracy on GlaS test sets with LayerCAM and PixelCAM with different stainings.
4 Conclusion
In this paper, PixelCAM is introduced for multi-task single-step WSOL method. It is based on a cost-effective FG/BG pixel-wise classifier in the pixel-feature space, allowing for spatial object localization. It leverages pixel-pseudo labels cues for localization. It aims at explicitly learning discriminant pixel-features and accurately separate FG/BG regions, promoting better ROI localization and image classification. Both tasks are optimized in parallel without incurring notable computational cost nor additional parameters. Our pixel classifier is versatile. It can easily be integrated into any CNN- or transform-based classifier architecture without modification. Results on two histology datasets showed that PixelCAM can improve WSOL baseline methods over both tasks, but also make them robust to OOD data. However, there is still room to improve our method by better optimizing both tasks while reducing their mutual negative impact. In addition, dealing with OOD in WSOL scenario is still an ongoing issue, and our method still requires more improvements despite its progress.
[…]
Part Appendix
[…]
Appendix A Related Work
WSOL has emerged as a low-cost training setup to localize objects while classifying an image content Choe et al. (2020); Murtaza et al. (2024); Zhou (2017). WSOL has also been extended to videos Belharbi et al. (2023, 2025). Moreover, there has been a recent focus on designing WSOL methods for histology image analysis Rony et al. (2023). This allows us to reduce the large annotation cost, in addition to building visually interpretable classifiers.
Single-step WSOL. Several approaches address the WSOL problem in a single step Choe et al. (2020); Murtaza et al. (2024); Rony et al. (2023) where a single model is trained to do both tasks, classification and localization. Usually, this is achieved by using a localization head followed by a spatial pooling layer to extract per-class probability. Early Ilse et al. (2018); Selvaraju et al. (2020); Zhou et al. (2016) but also recent WSOL works Wu et al. (2023); Zhu et al. (2024); Gao et al. (2021) follow this strategy. Although this has brought significant improvements into the field, these models still face several limitations. Since only image-class labels are used as supervision without any localization cues, CAM localization can have poor performance, more so when dealing with less salient objects such as in histology images. Recent work Rony et al. (2023) showed that without localization cues over histology images, these models often lead to highly unbalanced localization. This manifests either by under- or over-activations leading to high false negative/positive rates. In addition, this approach faces a great challenge in model selection as both tasks proceed in an asynchronous convergence Choe et al. (2020); Murtaza et al. (2024); Rony et al. (2023). For instance, the model selected for best localization often yields poor classification.
Two-step WSOL.This strategy has emerged as a new line of research. In particular, a dual-model approach is used in combination with the usage of localization cues under the form of pseudo-labels Murtaza et al. (2024); Rony et al. (2023). This provides direct guidance for the localization task, bypassing the issue of task convergence. Several works simply create a dedicated model per task: one for classification and another for localization Wei et al. (2021); Zhang et al. (2020); Zhao et al. (2023). This leads to better performance since tasks are divided between two large models. However, this is a cumbersome approach as it drastically increases the number of parameters and training cycles. Most importantly, the localization CAMs are completely disconnected from the classification decision, making this strategy unreliable and less interpretable. A parallel direction overcomes this issue by using a decoder to act as a localizer on top of the classifier Belharbi et al. (2022a, c); Murtaza et al. (2023), in a similar way to a U-Net architecture Ronneberger et al. (2015). This creates a direct relation between both tasks. However, the decoder has a limited learning capacity since it is tied to a frozen encoder at several layers via skip connections. This prevents the localizer from a better adaptation, and it limits its performance leading to a sub-optimal solution. In addition, it adds a significant number of parameters to the classifier.
WSOL vs OOD. In addition to these limitations, WSOL methods face challenges when dealing with domain shift which degrades the performance of both tasks Guichemerre et al. (2024). Such shift is common in histology due to variations in stains, objects’ structure, microscope type, and imaging centers. Further analysis shows that this issue could be rooted in the pixel features of the image encoder. Both classification and localization tasks depend heavily on these spatial features. Less flexible and poorly discriminant pixel features can lead to poor performance in both tasks since they both rely on these embeddings.
Appendix B Datasets
GlaS.
The GlaS dataset is used for the diagnosis of colon cancer. The dataset consists of 165 images from 16 Hematoxylin and Eosin (H&E) and includes labels at both pixel-level and image-level (benign or malign). The dataset consists of 67 images for training, 18 for validation, and 80 for testing. We use the same protocol as in (Rony et al., 2023; Guichemerre et al., 2024; Belharbi et al., 2022a). Similarly, we use 3 samples per class with full supervision in validation set for model selection for localization (B-LOC).
CAMELYON16.
A patch-based benchmark Rony et al. (2023) is extracted from the CAMELYON16 dataset that contains 399 Whole slide images categorized into two classes (normal and metastasic) for the detection of breast cancer metastases in H&E-stained tissue sections of sentinel lymph nodes. Patch extraction of size ${512\times 512}$ follows a protocol established by (Rony et al., 2023; Guichemerre et al., 2024; Belharbi et al., 2022a) to obtain patches with annotations at the image and pixel level. The dataset contains a total of 48870 images, including 24348 for training, 8850 for validation, and 15664 for testing. From the validation dataset, 6 examples per class are randomly selected to be fully supervised to perform model selection for localization (B-LOC) similarly to (Rony et al., 2023; Guichemerre et al., 2024; Belharbi et al., 2022a).
Appendix C WSOL Training Details
For pretraining a WSOL baseline method on the data, we use the same setup as in Guichemerre et al. (2024). The first part of the model training is defined using the backbone trained on ImageNet Krizhevsky et al. (2012). The training is done using SGD with a batch size of 32 Guichemerre et al. (2024). For the GlaS dataset, training is performed over 1000 epochs, and 20 epochs for CAMELYON16. A weight decay of $10^{-4}$ is also used. During training, images are resized to ${256\times 256}$, then randomly cropped to ${224\times 224}$. A hyperparameter search was conducted for the learning rate parameters among the values {0.0001, 0.001, 0.01}, and its decaying factor among {0.1, 0.4, 0.9} following Guichemerre et al. (2024). In the second phase of training PixelCAM, we use the CAMs generated by the previous method and continue the training using the same setup as defined previously. For the OOD scenario, the source model is trained using the same setup on a source dataset and evaluated on a target dataset.
Appendix D Class Separability Index
@@
In this section, we present the class separability index between foreground (FG) and background (BG) classes at pixel-feature level over test set of both datasets GlaS and CAMELYON16. To measure how well FG/BG classes are separated in the feature space of pixels, we resort to the class separability index Duda et al. (2000). The class separability, $J$, is based on the Within-class scatter matrix ($\bm{S}_{W}$) and the Between-class scatter matrix ($\bm{S}_{B}$) of pixel-features, defined as follows,
$\bm{S}_{W}=\sum_{i=1}^{c}\left[\sum_{j=1}^{n_{i}}(\bm{x}_{i,j}-\bm{m}_{i})(\bm{x}_{i,j}-\bm{m}_{i})^{\top}\right]\,,$ (4)
$\bm{S}_{B}=\sum_{i=1}^{c}n_{i}(\bm{m}_{i}-\bm{m})(\bm{m}_{i}-\bm{m})^{\top}\,,$ (5)
@@
where $c$ is the number of class (in our case ${c=2}$ for foreground and background). Note that this index $J$ is computed over a single image using the pixel-labels. Let $n_{i}\;(i=0,..,c)$ be the number of pixels in the $i$-th class. The feature vector $\bm{x}_{i,j}\in\mathbb{R}^{d}$ denotes the $j$-th pixel of the $i$-th class. It represents the vector ${\mathsf{F}_{p}}$ for the pixel $p$ in the main paper. The mean vector of the $i$-th class is given by $\bm{m_{i}}$, while $\bm{x}$ represents the mean vector computed over all feature vectors. The class separability denoted $J$ is defined as follows Duda et al. (2000),
$J=\frac{\operatorname{tr}(\bm{S}_{B})}{\operatorname{tr}(\bm{S}_{W})}\,.$ (6)
@@
In Table 4 and 5, we provide the average class separability over normal and cancer classes and the overview over the entire dataset. As we can observe, PixelCAM improves the class separability between FG and BG of WSOL baseline methods on both dataset except for GradCAM++ on CAMELYON16. This separability is explained by the high number of images with a low separability as illustrated in Fig 6. The improvement of PixelCAM in the OOD scenario compared to the WSOL baseline methods such as DeepMIL and LayerCAM on GlaS and CAMELYON16 can be attributed to better separability as shown in Figs 7 and 8. We provide examples of features separability at the pixel level in section K.
@@
Appendix E More Localization Performance Analysis
PxAP is the primary metric used to measure localization performance. However, following Rony et al. (2023), we include other pixel-wise performance measures that is true/false positives/negative rates. As shown previously, our method PixelCAM achieves better results in terms of PxAP performance. This improvement is observed in Tab. 6 with a higher number of true positive/negative rates compared to GradCAM++, LayerCAM, SAT, and NEGEV on the GlaS dataset. For the CAMELYON16 dataset, we observe that PixelCAM increases the true positive rate compared to standard methods while being competitive to NEGEV. However, we note a significant improvement when applying PixelCAM to the SAT method, with a considerable increase in true positives, thereby reducing the over-activation issue.